Fermi National Laboratory
 

Operation Summary for February 19, 2010 - February 22, 2010

 

The Integrated Luminosity for the period from 2/15/10 to 2/22/10 was 64.03 pb-1.

NuMI reported receiving 7.91E18 protons on target during this same period.

Friday Feb. 19

The day shift began with Operations monitoring the Tevatron (TeV) store 7619 (192.8E30) and the Recycler stash (144.1E10), with the Antiproton Source stacking (22.0E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beam.

During the day shift, a Linac low-energy RF station (LRF5) caused problems throughout the shift.

During the evening shift, Operations disabled beam to MTest at 6:05 PM. Operations terminated store 7619 at 7:05 PM and put the TeV into HEP shot setup. Operations established store 7620 at 8:30 PM with an Average Initial Luminosity (AIL) of 313.86E30.

Saturday Feb. 20

The midnight shift began with Operations monitoring store 7620 (174.44E30) and the stash (129.5E10), with Pbar stacking (15.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:20 AM. A MI RF station (MIRF4) tripped off and would not reset. Operators bypassed the station.

During the day shift, Operations terminated store 7620 at noon and put the TeV into shot setup. Operations established store 7621 at 1:15 PM with an AIL of 316.25E30.

During the evening shift, Operations disabled the MTest beam at 6:05 PM.

Sunday Feb. 21

The midnight shift began with Operations monitoring store 7621 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7621 at 2:05 AM. Operations prepared CDF for a collision hall access. Operations established beam to MTest at 4 AM. Operations put the TeV into shot setup at 5:15 AM and established store 7624 at 6:30 AM with an AIL of 310.74E30. MIRF9 tripped off due to a shorted screen power supply. A MI RF person investigated and determined the problem with MIRF9 was in the tunnel. He investigated the trouble with MIRF4 and began repairs on a faulty current sensor in its modulator.

During the day shift, MIRF4 was brought online by 9 AM. Operations reported that store 7625 had aborted at 10:55 AM due to an off site power glitch made the Master Substation glitch and caused two TeV RF stations to trip off. There was no quench. Operations put the TeV into shot setup and established store 7628 at 3:45 PM with an AIL of 177.24E30. A P150 transfer line vacuum burst caused some downtime while the vacuum recovered.

During the evening shift, the NuMI condensate tank system tripped off at 11:30 PM, which inhibited its beam.

Monday Feb. 22

The midnight shift began with Operations monitoring store 7628 and the stash, with Pbar stacking, with NuMI off, and with MiniBooNE taking beam.

During the midnight shift, a NuMI person brought a chiller back online; Operations resumed sending beam to NuMI at 1 AM. A Booster RF station (BRF1) tripped off at 3:25 AM; operators lowered the Booster beam intensities.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Cryo system work in between stores, and a Recycler access into MI31 and Booster RF work after the next store goes in.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/22/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory